OpenStudy (anonymous):

Can someone save me?? f(x) = x^2 - 2x and g(x) = x-4 find: h(x) = (f * g)(x)

OpenStudy (mathstudent55):

(fog)(x) = f(g(x)) That means use g(x) as input into f(x) f(x) = x^2, this is what goes into x of f(x) (fog)(x) = f(g(x)) = f(x-4) = (x - 4)^2 - 2(x - 4) Now simplify the right side

OpenStudy (anonymous):

x^2 - 2x +24?

OpenStudy (mathstudent55):

(fog)(x) = x^2 - 8x + 16 - 2x + 8 (fog)(x) = x^2 - 10x + 24

OpenStudy (anonymous):

where am i missing the 8x?

OpenStudy (mathstudent55):

How did you go from (x - 4)^2 - 2(x - 4) to x^2 - 2x +24?

OpenStudy (mathstudent55):

Remember that (x - 4)^2 = x^2 - 8x + 16

OpenStudy (anonymous):

oooooo right (x-4(x-4) I get it I get it